A student prepared a 3890ml solution of the solute Cs2CrO4 in water with a molarity of 2.45m.
What is the moles of solute in the solution?
What is the MM (g/mol) of the solute?
What is the mass of the solute required?
What is the density of the entire solution?
1
Expert's answer
2020-11-17T10:32:06-0500

Mole = molarity × volume = 2.45M × 3.890L = 9.5305moles

Molar mass = 2(133)+52+(16×4) = 382g/mol

Mass = molar mass × mole = 382×9.5305 = 3640.65g

Density = mass/volume = 3640.65g/3890mL = 0.9359g/mL